OpenStudy (anonymous):

To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended five sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write a system of two equations with the variables m and f in each to represent this problem do i use y=mx+b?

OpenStudy (anonymous):

Set up a system of equations. The membership fee is m and the fee per session is f. Your first equation could be: 25 = m + 5f Your second equation could be: 40 = m + 10f Solving this system will give you the solution.

OpenStudy (anonymous):

would i use addition,subtraction, or multiplication to solve the problem?

OpenStudy (anonymous):

Probably substitution is best. For example, isolate the 'm' variable in the first equation to obtain m = 25 - 5f. Then substitute this value of m into the second equation so your second equation becomes: 40 = (25 - 5f) + 10f. This will give you f, then you can solve for m using that value.

OpenStudy (anonymous):

ok i gte it thank you for the help